The height of a box is 10 inches. The length is three inches more than the width. Find the width if the volume is 540 cu inches

Answer by TutorDelphia(193)

Start with the formula
v=l*w*h
Write the info given in the problem
l=w+3
540=10*l*w
sub in w+3 for l
540=10(w+3)*w
Solve for w
540=10w(w+3)
540=10w^2+30w
0=10w^2+30w-540
0=10(w^2+3w-54)
0=10(w+9)(w-6)
w=-9 or 6
6 is the only reasonable answer
And to check 10*9*6=540